OpenStudy (anonymous):

What are the factors of 3x2 + 23x − 8?

OpenStudy (campbell_st):

(3x - 1)(x + 8)

OpenStudy (anonymous):

\[3x^2 + 24x - x -8 = 3x(x+8) - 1(x+8) = (x+8)(3x-1)\]